# Near Coincidences and Nilpotent Division Fields

### v1

@@ 3. Near Coincedences
In [29, p. 312], Serre explains that not only does $E$ gain good reduction over $\mathbb{Q}(\sqrt[12]{\Delta}),$ but also that ${\rm ord}_{p}(\Delta)\in\{0,2,3,4,6,8,9,10\}$. Thus, if $\mathfrak{p}$ is a prime above $p$ in $\mathbb{Q}(\sqrt[12]{\Delta})$, then $e(\mathfrak{p}/p)\in\{1,2,3,4,6\}.$
@@
Even with this in hand, for $p=5$ we will have to rely on the classification of rational points provided in [26]. To handle this case, we search $\operatorname{GL}_{2}(\mathbb{Z}/25\mathbb{Z})$ for groups that represent $(25,5)$ near coincidences. We then take the ones that are maximal with respect to containment (up to conjugation) and check if they have points. These groups are exactly the group with RSZB labels
@@
Using the data in [26], we see that there are non-cuspidal $\mathbb{Q}$-rational points on any of these curves and so there are no elliptic curves over $\mathbb{Q}$ with a $(25,5)$ near coincidence.
@@
Suppose towards a contradiction that this extension is not totally ramified and let $\mathfrak{P}$ be a prime above $\mathfrak{p}$. Then, $e(\mathfrak{P}/\mathfrak{p})$ is a proper divisor of $|\bar{\rho}_{E,p^{2}}(\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q}(\sqrt[12]{\Delta}))|$ which itself divides $2p(p^{2}-1).$ Thus
$e(\mathfrak{P}/\mathfrak{p})\leq p(p^{2}-1)\hbox{ and }e(\mathfrak{P}/p)=e(\mathfrak{P}/\mathfrak{p})e(\mathfrak{p}/p)\leq p(p^{2}-1)\cdot 6<p^{2}(p^{2}-1).$
@@
Let $E/\mathbb{Q}$ be an elliptic curve and let $p\geq 7$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of the normalizer of a non-split Cartan subgroup $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$ Then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
@@
[…]
The last case that remains here is the case that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in an exceptional group. These curves are completely classified and again using [26], we see that in this case $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$ cannot occur.
@@
To start this section, we push a little further in the cases where $p=2$ and $3$ In both of these cases, we can have near coincidence between the $p^{2}$- and $p$-division fields, but what about the $p^{3}$- and $p^{2}$-division fields?
So we search for groups that represent $(8,4)$ and $(27,9)$ coincidences. In the first case, we find that the maximal groups that represent an $(8,4)$ coincience all have genus 1 or higher. Using the data in [27], we know that this means that there is no elliptic curve without complex multiplication that has these images, and we have already completely dealt with the CM case.
@@
Next we prove the case of Theorem 1.3 when $n\geq 2$ by inductcion.
@@
Suppose that $E/\mathbb{Q}$ is an elliptic curve and $p\in\mathbb{Z}$ a prime greater than 2 such that $p^{k}$ divides $[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ for some $k\in\{1,2,3,4\}$ and $n\geq 1.$ Then, $p^{k}$ divides $[\mathbb{Q}(E[p^{n+2}]):\mathbb{Q}(E[p^{n+1}])]$
@@
Assume that $n\geq 1$ and $p^{k}$ divides $[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ for some $k\in\{1,2,3,4\}$ This means that the set
@@
$\tilde{S}=\{A\in\operatorname{Im}\bar{\rho}_{E,p^{n+2}}\colon A\equiv I\bmod{p^{n+1}}\}$
Notice that if $I+p^{n}X\in S$ for some $X\in M_{2}(\mathbb{Z}/p\mathbb{Z})$, then there is a $\sigma_{0}\in\operatorname{Gal}(\mathbb{Q}(E[p^{n+1}]/\mathbb{Q})$ such that $\bar{\rho}_{E,p^{n+1}}(\sigma_{0})=I+p^{n}X$. Further, there must be a $\sigma\in\operatorname{Gal}(\mathbb{Q}(E[p^{n+2}]/\mathbb{Q})$ such that $\sigma\big{|}_{\mathbb{Q}(E[p^{n+1}])}=\sigma_{0}$. In this case, since $\bar{\rho}_{E,p^{n+1}}(\sigma_{0})\equiv I+p^{n}X$ we have that
@@
for some $X\in M_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ such that $\tilde{X}\equiv X\bmod p$. Then it must be that
@@
$\equiv I+p^{n+1}\tilde{X}\bmod p^{n+2}.$
Thus $I+p^{n+1}\tilde{X}\in\tilde{S}$ and it must be that $|\tilde{S}|\geq|S|$. This forces $\mathbb{Q}(E[p^{n+2}])/\mathbb{Q}(E[p^{n+1}])$ to have degree at least $|S|\geq p^{k}$ and so
$p^{k}\mid[\mathbb{Q}(E[p^{n+2}]):\mathbb{Q}(E[p^{n+1}]).$
@@
Notice that if $p\geq 3$, then the statement that $p^{2}\mid[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ is equivalent to the fact that $\mathbb{Q}(E[p^{n+1}])\neq\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}).$ This is because the field extension $\mathbb{Q}(E[p^{n+1}])/\mathbb{Q}(E[p^{n}])$ is a Galois extension whose Galois group is isomorphic to a subgroup of the additive group of $2\times 2$ matrices with entries in $\mathbb{Z}/p\mathbb{Z}$, $M_{2}(\mathbb{Z}/p\mathbb{Z})$. The group $M_{2}(\mathbb{Z}/p\mathbb{Z})$ has order $p^{4}$ and so a priori $[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]=p^{k}$ for some $k\in\{0,1,2,3,4\}.$ We omit the case when $k=0$ in Proposition 3.9 since it is uninteresting.
@@
$[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]=[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}})][\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}):\mathbb{Q}(E[p^{n}])].$
But, by Theorem 1.1, in this case $\zeta_{p^{n+1}}\not\in\mathbb{Q}(E[p^{n}])$ and so $[\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}):\mathbb{Q}(E[p^{n}])]=p$. Bringing it all together we see that
@@
Suppose that $p=2$ and $E/\mathbb{Q}$ is an elliptic curve such that $p^{k}$ divides $[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ for some $k\in\{1,2,3,4\}$ and $n\geq 2.$ Then, $p^{k}$ divides $[\mathbb{Q}(E[p^{n+2}]):\mathbb{Q}(E[p^{n+1}])]$
@@ 4.1. Nilpotent Groups
A group $G$ is called nilpotent if it has a normal central series.
@@
[16, Theorem 1.26] Let $G$ be a finite non-trivial group. The following are equivalent:
@@
[6, Theorem 5.7] Nilpotencty is closed under subgroups, quotients, and direct products.
@@
If $|Z(G)|=1$, no term in an ascending central series can have order larger than $1$. ∎
@@ 4.2. Classification of nilpotent division fields of prime level
Step one in the process of determining when an elliptic curve $E/\mathbb{Q}$ can have a nilpotent $n$-division field, is determining when the $p$-division fields can be nilpotent extensions of $\mathbb{Q}$. Proposition 4.4 tells us that if $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent, then $\mathbb{Q}(E[d])/\mathbb{Q}$ is nilpotent for all $d\mid n$. Moreover, if $n=p_{1}^{a_{1}}\cdots p_{k}^{a_{k}}$ is the prime factorization of $n$, then $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent if and only if $\mathbb{Q}(E[p_{i}^{a_{i}}])/\mathbb{Q}$ is nilpotent for all $i$. To see why this is true, one only needs to recall the Galois correspondence as well as the fact that nilpotency is perserved under subgroups and quotients.
@@
In order to keep this statement as clean as possible, we will be adopting the convention that $(\mathbb{Z}/2\mathbb{Z})^{2}$ is a dihedral group. Consider the group
$G:=\left\langle\begin{pmatrix}2&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\rangle\subseteq\operatorname{GL}_{2}(\mathbb{Z}/3\mathbb{Z}).$
The group $G$ is a Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/3\mathbb{Z})$ and it is isomorphic to $D_{4}$. Its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{2}$ and so to include this case in case (2) of Proposition 2.1, we consider $D_{2}$ to be the Klein 4-group.
Since $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is nilpotent if and only if its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is nilpotent, we focus on when the latter can occur. First note that if $p\mid|G|$, and contains $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$, then $G$ cannot be nilpotent since $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is not nilpotent.
Next, suppose that $G$ is an admissible nilpotent group such that $p\mid|G|$, and $G$ is contained in the Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. Then, using the fact that $p$ divides the size of $G$, it follows that $G$ cannot be contained inside the split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. Conjugating $G$ if necessary, we may assume $G$ is contained in the set of upper triangular matrices and that
$A=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$
is in $G$. This forces the $p$-Sylow subgroup of $G$ to be exactly the group generated by $A$. Then, since we assumed that $G$ was nilpotent, we have from part (3) of Theorem 4.3 every element in $G$ must commute with $A.$ Let
@@
be an arbitrary element of $G$. Notice we can assume that $B$ is of this form since $G$ is contained in a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ and thus is conjugate to a subgroup of the upper triangular matrices. Next we compute
@@
$G=\left\langle\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\rangle\subseteq\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z}).$
Next we note that all of the groups $G$, that fall into part (3) of the case when $p\nmid|G|$ are not nilpotent by Proposition 4.5, since each of $A_{4}$, $S_{4}$ and $A_{5}$ have trivial centers.
So the last remaining case to deal with is part when $G$ is contained in the normalizer of a Cartan subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$ In this case, Example 4.6 gives that $G$ is nilpotent if and only if its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ has order a power of $2$. The previous discussion can be summarized in the following proposition.
Proposition 4.9.
Let $p$ be a prime and let $G$ be an admissible subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ such that $G$ is nilpotent. Then, $G$ is abelian or the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is isomorphic to $D_{2^{k}}$ for some $k\geq 2$. Further, if $p$ is odd, then $G$ is either contained in the normalizer of a split or non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$.

### v3

@@ 3. Near coincidences
In [36, p. 312], Serre explains that not only does $E$ gain good reduction over $\mathbb{Q}(\sqrt[12]{\Delta}),$ but also that ${\rm ord}_{p}(\Delta)\in\{0,2,3,4,6,8,9,10\}$. Thus, if $\mathfrak{p}$ is a prime above $p$ in $\mathbb{Q}(\sqrt[12]{\Delta})$, then $e(\mathfrak{p}/p)\in\{1,2,3,4,6\}.$
Even with this in hand, for $p=5$ we will have to rely on the classification of rational points provided in [33].
Proposition 3.7.
Let $E/\mathbb{Q}$ with the property that $\mathrm{im}\,\bar{\rho}_{E,5}$ is contained in the normalizer of the non-split Cartan subgroup mod $5$. Then $\mathbb{Q}(E[25])\neq\mathbb{Q}(E[5],\zeta_{25})$.
Proof.
@@
We search $\operatorname{GL}_{2}(\mathbb{Z}/25\mathbb{Z})$ for groups that represent $(25,5)$ near coincidences. We then take the ones that are maximal with respect to containment (up to conjugation) and check if they have points. These groups are exactly the group with RSZB labels
@@
Using the data in [33], we see that there are no non-cuspidal $\mathbb{Q}$-rational points on any of these curves and so there are no elliptic curves over $\mathbb{Q}$ with a $(25,5)$ near coincidence. ∎
We now turn to the case that $p\geq 7$.
@@
Let $E/\mathbb{Q}$ be an elliptic curve and let $p\geq 7$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is a subgroup of the normalizer of a non-split Cartan subgroup $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$ Further, let $\Delta$ be the discriminant of $E$ and suppose that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$. Let $\mathfrak{p}$ be a prime over $p$ in $\mathbb{Q}(\sqrt[12]{\Delta}).$ The extension $\mathbb{Q}(\sqrt[12]{\Delta},E[p^{2}])/\mathbb{Q}(\sqrt[12]{\Delta})$ is a degree $2p(p^{2}-1)$ extension that is totally ramified at $\mathfrak{p}$.
Proof.
Suppose towards a contradiction that there is a prime $\mathfrak{P}$ above $\mathfrak{p}$ such that $e(\mathfrak{P}/\mathfrak{p})$ is a proper divisor of $2p(p^{2}-1)$. Since $\operatorname{Im}\bar{\rho}_{E,p}$ is a subgroup of the normalizer of a non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$, we know that $\mathbb{Q}(E[p]):\mathbb{Q}]$ is a divisor of $2(p^{2}-1)$. Further, because we assumed that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$, we know that $\mathbb{Q}(E[p^{2}]):\mathbb{Q}]$ is a divisor of $2p(p^{2}-1)$. Thus $\mathbb{Q}(\sqrt[12]{\Delta},E[p^{2}])/\mathbb{Q}(\sqrt[12]{\Delta})$ has degree dividing $2p(p^{2}-1)$. Now, $e(\mathfrak{P}/\mathfrak{p})$ is a proper divisor of $|\bar{\rho}_{E,p^{2}}(\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q}(\sqrt[12]{\Delta}))|$ which itself divides $2p(p^{2}-1).$ Thus
@@
We now explain how the previous proposition rules out the possibility of a $(p^{2},p)$ near coincidence.
Proposition 3.9.
Let $E/\mathbb{Q}$ be an elliptic curve and let $p\geq 7$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is a subgroup of the normalizer of a non-split Cartan subgroup $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z}).$ Then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
@@
Suppose towards a contradiction that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$. By the previous proposition, we know that $\mathbb{Q}(\sqrt[12]{\Delta},E[p^{2}])/\mathbb{Q}(\sqrt[12]{\Delta})$ is totally ramified at $\mathfrak{p}$. The assumption that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$ implies that $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ has order dividing $2p(p^{2}-1)$. On the other hand, Proposition 3.8 implies that
$|\bar{\rho}_{E,p^{2}}(\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q}(\sqrt[12]{\Delta})))|=2p(p^{2}-1).$
It follows from this that
$\operatorname{Im}\bar{\rho}_{E,p^{2}}=\bar{\rho}_{E,p^{2}}(\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q}(\sqrt[12]{\Delta})))$
and hence
$\operatorname{Gal}(\mathbb{Q}(\sqrt[12]{\Delta},E[p^{2}])/\mathbb{Q}(\sqrt[12]{\Delta}))\simeq\operatorname{Im}\bar{\rho}_{E,p^{2}}.$
However, a Galois extension that is totally ramified at a prime over $p$ must have a Galois group which is an extension of a finite $p$-group (the wild inertia group) by a finite cyclic group of order coprime to $p$ (the tame inertia group). This means that for any prime $q\neq p$, the $q$-Sylow subgroup of $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ must be cyclic. In particular, the Sylow $2$-subgroup of $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ must be cyclic. However, by [16], $\mathbb{Q}(E[p])/\mathbb{Q}$ must be nonabelian, and so the projectivization of $\operatorname{Im}\bar{\rho}_{E,p}$ must have a dihedral Sylow $2$-subgroup. This is a contradiction. ∎
The last case that remains here is the case that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in an exceptional group (i.e. the image in $\operatorname{PGL}_{2}(\mathbb{F}_{p})$ falls into case 3 of Proposition 2.1).
@@
Let $E/\mathbb{Q}$ be an elliptic curve and let $p\geq 5$ be a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in an exceptional subgroup. Then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
@@
Serre showed [37, Section 8.4, Lemme 18] that $\operatorname{Im}\bar{\rho}_{E,p}$ can only be an exceptional subgroup if image in $\operatorname{PGL}_{2}(\mathbb{F}_{p})$ is isomorphic to $S_{4}$, $p\leq 13$ and $p\equiv 3\text{ or }5\pmod{8}$.
The elliptic curves with $\operatorname{Im}\bar{\rho}_{E,13}$ contained in an exceptional subgroup were determined in [2] and in [33] it was shown that for each such elliptic curve, $\operatorname{Im}\bar{\rho}_{E,13^{2}}$ contains all matrices $\equiv I\pmod{13}$, which implies that $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
For $p=11$, the elliptic curves with $\operatorname{Im}\bar{\rho}_{E,11}$ contained an exceptional subgroup were determined by Ligozat [26, Proposition II.4.4.8.1] and the only possibility is $j(E)=0$. For such an elliptic curve, we cannot have $\mathbb{Q}(E[11^{2}])=\mathbb{Q}(E[11],\zeta_{11^{2}})$ by Proposition 2.7.
For $p=5$, if $\operatorname{Im}\bar{\rho}_{E,5}$ is contained in an exceptional subgroup, then it either equals an exceptional subgroup, or is contained in a Borel or the normalizer of a split Cartan. The latter two cases are impossible by Proposition 3.4 and Proposition 3.5. If the image is the exceptional subgroup mod $5$, then a group theory computation with Magma shows that the mod $25$ image of Galois has RSZB label 25.625.36.1, which is shown to be impossible in [33, Subsection 8.6]. ∎
@@
To start this section, we push a little further in the cases where $p=2$ and $3.$ These searches were carried out using a Magma script, NearCoin.m, which can be found in [8]. In both of these cases, we can have near coincidence between the $p^{2}$- and $p$-division fields, but what about the $p^{3}$- and $p^{2}$-division fields?
So we search for groups that represent $(8,4)$ and $(27,9)$ coincidences. In the first case, we find that the maximal groups that represent an $(8,4)$ coincidence all have genus 1 or higher. Using the data in [34], we know that this means that there is no elliptic curve without complex multiplication that has these images, and we have already completely dealt with the CM case.
@@
Next we prove the case of Theorem 1.13 when $n\geq 2$ by induction.
@@
Suppose that $E/\mathbb{Q}$ is an elliptic curve and $p>2$ is a prime such that $p^{k}$ divides $\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ for some $k\in\{1,2,3,4\}$ and $n\geq 1.$ Then, $p^{k}$ divides $\mathbb{Q}(E[p^{n+2}]):\mathbb{Q}(E[p^{n+1}])]$
@@
Assume that $n\geq 1$ and $p^{k}$ divides $\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ for some $k\in\{1,2,3,4\}$ This means that the set
@@
$\tilde{S}=\{A\in\operatorname{Im}\bar{\rho}_{E,p^{n+2}}\colon A\equiv I\bmod{p^{n+1}}\}.$
[…]
We can represent an element of $S$ in the form $I+p^{n}X$ with $X$ in the additive group $M_{2}(\mathbb{Z}/p\mathbb{Z})$ of $2\times 2$ matrices. Define $\phi:S\to\tilde{S}$ by $\phi(I+p^{n}X)=I+p^{n+1}X$. It is straightforward to see that this formula defines an injective homomorphism, but it is not immediately clear that if $I+p^{n}X\in S$, then $I+p^{n+1}X\in\tilde{S}$. We now justify this.
If $I+p^{n}X\in S$ for some $X\in M_{2}(\mathbb{Z}/p\mathbb{Z})$, then there is a $\sigma_{0}\in\operatorname{Gal}(\mathbb{Q}(E[p^{n+1}]/\mathbb{Q}))$ such that $\bar{\rho}_{E,p^{n+1}}(\sigma_{0})=I+p^{n}X$. Further, there must be a $\sigma\in\operatorname{Gal}(\mathbb{Q}(E[p^{n+2}])/\mathbb{Q})$ such that $\sigma\big|_{\mathbb{Q}(E[p^{n+1}])}=\sigma_{0}$. In this case, since $\bar{\rho}_{E,p^{n+1}}(\sigma_{0})\equiv I+p^{n}X$ we have that
@@
for some $\tilde{X}\in M_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ such that $\tilde{X}\equiv X\bmod p$. Then it must be that
@@
$\equiv I+p^{n+1}\tilde{X}\bmod p^{n+2}$
$\equiv I+p^{n+1}X\bmod p^{n+2}.$
Thus $I+p^{n+1}X\in\tilde{S}$.
[…]
$p^{k}\mid[\mathbb{Q}(E[p^{n+2}]):\mathbb{Q}(E[p^{n+1}])].$
@@
Notice that if $p\geq 3$, then the statement that $p^{2}\mid[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ is equivalent to the fact that $\mathbb{Q}(E[p^{n+1}])\neq\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}).$ This is because the field extension $\mathbb{Q}(E[p^{n+1}])/\mathbb{Q}(E[p^{n}])$ is a Galois extension whose Galois group is isomorphic to a subgroup of the additive group $M_{2}(\mathbb{Z}/p\mathbb{Z})$ of $2\times 2$ matrices with entries in $\mathbb{Z}/p\mathbb{Z}$. The group $M_{2}(\mathbb{Z}/p\mathbb{Z})$ has order $p^{4}$ and so a priori $\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]=p^{k}$ for some $k\in\{0,1,2,3,4\}.$ We omit the case when $k=0$ in Proposition 3.13 since it is uninteresting.
@@
$\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]=[\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}})][\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}):\mathbb{Q}(E[p^{n}])].$
But, by Theorem 1.12, in this case $\zeta_{p^{n+1}}\not\in\mathbb{Q}(E[p^{n}])$ and so $\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}):\mathbb{Q}(E[p^{n}])]=p$. Bringing it all together we see that
@@
Suppose that $p=2$ and $E/\mathbb{Q}$ is an elliptic curve such that $p^{k}$ divides $\mathbb{Q}(E[p^{n+1}]):\mathbb{Q}(E[p^{n}])]$ for some $k\in\{1,2,3,4\}$ and $n\geq 2.$ Then, $p^{k}$ divides $\mathbb{Q}(E[p^{n+2}]):\mathbb{Q}(E[p^{n+1}])]$
@@ 4.1. Nilpotent Groups
A group $G$ is called nilpotent if it has a central series.
@@
[19, Theorem 1.26] Let $G$ be a finite non-trivial group. The following are equivalent:
@@
[5, Theorem 5.7] Nilpotency is closed under subgroups, quotients, and direct products.
@@
Suppose $G$ is a finite nontrivial group that is nilpotent. Then, by Theorem 4.3 part 3, $G$ is the direct products of its $p$-Sylow subgroups. A classical result in group theory is that $p$-group have nontrivial centers and so $G$ must have a nontrivial center. ∎
@@
Before proceeding, we remind the reader of the following definition:
Definition 4.8.
Let $K/k$ be a Galois extension of fields. We say that $K/k$ is a nilpotent extension if $\operatorname{Gal}({K}/k)$ is a nilpotent group. When the base field $k$ is clear from context, we will just say that $K$ is a nilpotent field for brevity.
@@ 4.2. Classification of nilpotent division fields of prime level
Step one in the process of determining when an elliptic curve $E/\mathbb{Q}$ can have a nilpotent $n$-division field, is determining when the $p$-division fields can be nilpotent extensions of $\mathbb{Q}$. Proposition 4.4 tells us that if $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent, then $\mathbb{Q}(E[d])/\mathbb{Q}$ is nilpotent for all $d\mid n$. Moreover, if $n=p_{1}^{a_{1}}\cdots p_{k}^{a_{k}}$ is the prime factorization of $n$, then $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent if and only if $\mathbb{Q}(E[p_{i}^{a_{i}}])/\mathbb{Q}$ is nilpotent for all $i$. To see why this is true, one only needs to recall the Galois correspondence as well as the fact that nilpotency is preserved under subgroups, quotients, and direct products.
@@
If $G$ is a subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$, then $G$ is nilpotent if and only if its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is nilpotent. The reason is as follows. If $G$ is nilpotent, then its image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is a quotient of $G$ and is hence nilpotent. Conversely, if $G\leq\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ and the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is nilpotent, then the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is the quotient $G/N$, where $N$ is the set of scalar multiples of the identity in $G$. This subgroup $N\leq Z(G)$ and by Theorem 5.13 of [5], it follows that $G$ is nilpotent, since $N$ can be extended into a central series.
The following result gives a classification of when an admissible subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is nilpotent.
Proposition 4.10.
Let $p$ be a prime and let $G$ be an admissible subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. Then $G$ is nilpotent if and only if $G$ is abelian, or the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is isomorphic to $D_{2^{k}}$ for some $k\geq 2$. Further, if $p$ is odd, then $G$ is either contained in the normalizer of a split or non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$.
Proof.
If $G$ is abelian, then it must be nilpotent. Remark 4.9 shows that if the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is isomorphic to $D_{2^{k}}$ (a $2$-group) then $G$ is nilpotent.
Now we assume that $G$ is nilpotent and consider the cases based on Proposition 2.1. First, if $G$ contains $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$, then $G$ cannot be nilpotent since $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$ is not nilpotent.
Next, suppose that $G$ is an admissible nilpotent group such that $p\mid|G|$, and $G$ is contained in a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. Conjugating $G$ if necessary, we may assume $G$ is contained in the set of upper triangular matrices. The set of upper triangular matrices contains a unique subgroup of order $p$, namely the cyclic group generated by
$A=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$.
Since we assumed that $G$ was nilpotent, we have from part (3) of Theorem 4.3 every element in $G$ must commute with $A.$ Let
@@
be an arbitrary element of $G$. Next we compute
@@
$G=\left\langle\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\rangle\subseteq\operatorname{GL}_{2}(\mathbb{Z}/2\mathbb{Z}),$
which is abelian.
We now consider case (1) of Proposition 2.1. In this case, $G$ is contained in a Cartan subgroup and is hence abelian.
In case (2) of Proposition 2.1, the projective image is dihedral and by Example 4.6, a dihedral group is nilpotent if and only if its order is a power of $2$.
In case (3) of Proposition 2.1 the projective image is isomorphic to $A_{4}$, $S_{4}$ or $A_{5}$. If $G$ were nilpotent this would imply that one of $A_{4}$, $S_{4}$ or $A_{5}$ is nilpotent. Proposition 4.5 shows that none of these are nilpotent, since all have trivial center. ∎